# m2mKD: Module-to-Module Knowledge Distillation for Modular Transformers

### v1

@@ Abstract
Modular neural architectures are gaining increasing attention due to their powerful capability for generalization and sample-efficient adaptation to new domains. However, training modular models, particularly in the early stages, poses challenges due to the optimization difficulties arising from their intrinsic sparse connectivity. Leveraging the knowledge from monolithic models, using techniques such as knowledge distillation, is likely to facilitate the training of modular models and enable them to integrate knowledge from multiple models pretrained on diverse sources. Nevertheless, conventional knowledge distillation approaches are not tailored to modular models and can fail when directly applied due to the unique architectures and the enormous number of parameters involved. Motivated by these challenges, we propose a general module-to-module knowledge distillation (m2mKD) method for transferring knowledge between modules. Our approach involves teacher modules split from a pretrained monolithic model, and student modules of a modular model. m2mKD separately combines these modules with a shared meta model and encourages the student module to mimic the behaviour of the teacher module. We evaluate the effectiveness of m2mKD on two distinct modular neural architectures: Neural Attentive Circuits (NACs) and Vision Mixture-of-Experts (V-MoE). By applying m2mKD to NACs, we achieve significant improvements in IID accuracy on Tiny-ImageNet (up to 5.6%) and OOD robustness on Tiny-ImageNet-R (up to 4.2%). On average, we observe a 1% gain in both ImageNet and ImageNet-R. The V-MoE-Base model trained using m2mKD also achieves 3.5% higher accuracy than end-to-end training on ImageNet. The experimental results demonstrate that our method offers a promising solution for connecting modular networks with pretrained monolithic models. Code is available at https://github.com/kamanphoebe/m2mKD.
@@ 1 Introduction
Despite the remarkable success of large monolithic models in various domains, such as computer vision and natural language processing, concerns have arisen regarding their limited generalization ability and increasing computational costs. In the meantime, there has been a growing interest in modular models, which promise to mitigate the drawbacks of monolithic models. In contrast to monolithic models with fixed computational graphs and parameters, modular neural architectures adapt their parameters to the input, offering favorable properties that are absent in static monolithic models (Han et al.,, 2022). Unlike monolithic models that collectively optimize parameters, modular models consist of independent modules, and each module can be updated locally without affecting other network parts. These modules are trained to specialize in specific tasks. During inference, only the relevant modules are activated based on the input, even for out-of-distribution samples, thereby enhancing generalization performance (Pfeiffer et al.,, 2023). For instance, DEMix Layers (Gururangan et al.,, 2022) jointly represent COVID-19-related data using medical and news modules learned from training data. Besides, the conditional calculation in modular models improves computational efficiency. A typical modular architecture, Mixture-of-Experts (MoE) (Shazeer et al.,, 2017), dramatically scales up model capacity while maintaining similar computational requirements as the original model (Jia et al.,, 2021; Lepikhin et al.,, 2021).
@@
Although modular architectures surpass monolithic models in terms of out-of-distribution (OOD) robustness and computational efficiency, there is still vast room for improvement in their algorithms and implementations. Training modular models presents challenges due to optimization difficulties arising from sparse interactions. While recent works (Zoph et al.,, 2022; Mustafa et al.,, 2022; Nie et al.,, 2022) have investigated the training instability of modular models, thorough studies in this area still need to be completed. On the engineering side, several implementations of MoE exist, but many of them (He et al.,, 2022; Lepikhin et al.,, 2021) do not consider dynamic workload allocation for experts or support acceleration techniques like mixed precision training. Although adaptive parallelism (Hwang et al.,, 2023) has addressed the dynamic nature of MoEs, training MoEs remains considerably slower than their monolithic counterparts due to unavoidable communication overheads.
@@
To alleviate the optimization issue, employing a pretrained monolithic model to guide the training of modular models shows potential. Knowledge distillation (KD) (Hinton et al.,, 2015) is an existing technique that transfers knowledge from a pretrained teacher model to a smaller student model. KD has proven effective in the context of monolithic models. However, directly applying conventional KD approaches to modular models is computationally expensive due to their large model sizes. Using a monolithic model as the teacher for a larger modular model may even harm performance (see Table 6). Furthermore, monolithic models trained by regular methods may not be the optimal choice for teachers.
@@
Inspired by the divide-and-conquer training mechanism of Deep Incubation (Ni et al.,, 2023), we introduce module-to-module knowledge distillation (m2mKD) to transfer knowledge between sub-modules of the monolithic teacher and the modular student. As illustrated in Figure 1, we first adopt Deep Incubation, a modular training method, to incubate the teacher modules using a small meta-model. Next, we encourage the student module to imitate the behaviour of the teacher module. Finally, the distilled student modules are used to initialize the modular student model. By adapting the distillation process at the distributed module level, m2mKD significantly reduces the capacity requirement of the teacher model and enables independent training, making distillation less computationally expensive. The cost savings become more noticeable when the modular student is large or when multiple teachers are involved, such as in ensemble learning. Moreover, the teacher model is trained in a modular way, which may benefit teaching a modular student. Note that our m2mKD algorithm does not impose any restrictions on the architecture of both the teacher and student models. We evaluate the performance of m2mKD on both NACs and V-MoE. Experiments show that a NAC model trained using m2mKD improves IID accuracy by approximately 5% on Tiny-ImageNet and enhances out-of-distribution (OOD) robustness by about 4% on Tiny-ImageNet-R. Additionally, there is an average gain of approximately 1% on ImageNet (Deng et al.,, 2009) and ImageNet-R (Hendrycks et al., 2021a, ). The experimental results for V-MoE models indicate that m2mKD also works in the case of a small teacher.
@@
Our contributions are as follows: (1) To the best of our knowledge, we are the first to investigate knowledge distillation for modular models. (2) We demonstrate the challenges associated with distilling modular models and introduce a tailored algorithm to address these challenges. (3) Our approach can be seen as a promising framework for transforming a monolithic model into a modular model with arbitrary neural architecture. Notably, this transformation is developed in a modular-style manner. (4) The proposed method is capable of handling irregular distillation scenarios. It has the potential to work not only for the monolithic-to-modular case but also for the monolithic-to-monolithic case. (5) We verify the feasibility of using Deep Incubation for modular models.
@@ 2 Related Work
Knowledge Distillation. Knowledge distillation is a model compression technique that transfers knowledge from a large teacher network to a smaller student model (Hinton et al.,, 2015). Common knowledge distillation approaches involve mimicking the teacher model’s softened output (Hinton et al.,, 2015; Kim et al.,, 2018; Ba and Caruana,, 2014). In scenarios with multiple teachers, the outputs of all teacher models can be averaged (Hinton et al.,, 2015), or other methods can be used (Fukuda et al.,, 2017; Kwon et al.,, 2020; Yuan et al.,, 2021). Traditional knowledge distillation methods often struggle when a significant capacity gap exists between the student and the teacher models. Recent approaches have attempted to address this gap by using teacher assistants (Mirzadeh et al.,, 2019; Son et al.,, 2021). In addition, some existing works conduct distillation at the module level instead of distilling entire models Xu et al., (2020); Zhao et al., (2022); Liang et al., (2023). While previous works primarily focus on compressing model sizes within the context of monolithic models, we propose a module-to-module distillation technique to enhance the performance of modular architectures, especially in unconventional distillation scenarios. Particularly, the distillation of modules is mutually independent and can be executed in a distributed manner.
@@
Modularization. Modular deep learning involves decomposing neural architectures into independent and parameter-efficient modules, where samples are conditionally routed to subsets of modules, and their outputs are aggregated. Modularization has been widely applied in various areas such as transfer learning (Li and Liang,, 2021; Houlsby et al.,, 2019; Platanios et al.,, 2018; Hu et al.,, 2021), modular training (Ni et al.,, 2023), and scaling up model size (Jia et al.,, 2021; Lepikhin et al.,, 2021).
@@
In transfer learning, a common approach is to use pre-trained models as modules in an assembled model for new tasks. The assembled model adapts to new data by adjusting some modules or adding new modules to improve performance in new scenarios. For example, Brown et al. (Brown et al.,, 2020) added a prompt module to the input of a pre-trained model, and Rebuffi et al. (Rebuffi et al.,, 2018) introduced adapter modules into the model architecture. In modular training, Deep Incubation (Ni et al.,, 2023) incubates modules in individual nodes, which avoids communication overhead between nodes and accelerates convergence. Conditional computation allows for scaling up model size while maintaining inference complexity. V-MoE (Jia et al.,, 2021) scales up vision models with only half the computation required during inference. The MoE framework introduces experts to achieve modularization specifically at the FFN layer. Other modular model architectures include NACs (Weiss et al.,, 2022) and (Shen et al.,, 2023; Goyal et al.,, 2021).
@@
Monolithic to Modular. There are several efforts to convert monolithic models into modular architectures. MoEfication (Zhang et al., acl, ) directly splits the FFN layers of a monolithic model into multiple experts to form MoE layers, while Sparse Upcycling (Komatsuzaki et al.,, 2022) copies the MLP parameters to the corresponding experts in the MoE layers. In contrast to these works, our proposed m2mKD does not make assumptions about the model architecture and can be applied to any model.
@@ 3 Method
In this section, we first review the essential aspects of Deep Incubation, as our work heavily relies on it. Subsequently, we introduce our proposed method. For convenience, we provide a table (Table 8) that lists the notations commonly used in this paper.
Figure 2: Overview of Deep Incubation. Initially, a small model is pre-trained as a meta model. Subsequently, the module replaces the corresponding layer in the meta-model, and only the parameters of the module are updated during incubation. Finally, the incubated modules are assembled and fine-tuned to obtain the final model.
@@ 3.1 Preliminary: Deep Incubation
Deep Incubation (Ni et al.,, 2023) employs a divide-and-conquer strategy to train large models modularly, as illustrated in Figure 2. The process consists of three stages: meta-model pre-training, module incubation, and assembly. In the initial stage, a small model is pre-trained as a meta model using an end-to-end training approach. Subsequently, each module replaces the corresponding layer in the meta-model, and only the module parameters are updated during the incubation process. The incubation of modules is mutually independent and can be completed in a distributed manner. Finally, the incubated modules are assembled and fine-tuned to obtain the final model. It is worth noting that the number of layers in the meta model corresponds to the number of modules incubated during the incubation stage.
@@ 3.2 Pipeline
Our method follows a pipeline consisting of three steps (Figure 3): preparation, module-to-module knowledge distillation (m2mKD), and end-to-end (E2E) training. More details are presented below.
@@
As proposed by (Ni et al.,, 2023), the initial step involves training the meta model in an end-to-end fashion to incubate the $L$ sub-modules $T_{1},T_{2},...,T_{L}$. Subsequently, these resulting sub-modules are reassembled to form the teacher model $\mathcal{T}$, which undergoes fine-tuning. While we follow the same pipeline described in the original paper, our focus is on the sub-modules themselves rather than the entire model. We opt for the Incubation approach instead of merely separating a pre-trained model because we claim that the module incubation phase imparts additional knowledge to the sub-modules. This allows them to learn how to function as individual modules rather than being incomplete fragments of a whole model.
Module-to-module Knowledge Distillation. Once the fine-tuning of the assembled model $\mathcal{T}$ is complete, the sub-modules, or we call “teacher modules” hereafter, are ready for running m2mKD. Unlike traditional knowledge distillation approaches, m2mKD aims to transfer knowledge between modules rather than entire models. Similar to the module incubation proposed by (Ni et al.,, 2023), we separately link the teacher and student module to the meta model. By comparing the outputs of the two resulting hybrid models, the student modules are encouraged to mimic the behaviour of the corresponding teacher modules. Previous research by Yang et al. (Yang et al.,, 2022) demonstrates that blocks located at similar depths in different networks can be considered functionally equivalent. This insight suggests that neural networks learn similar patterns at similar network stages. Exploiting this insight, we assign teacher modules to students at the same depth, resulting in $L$ teaching pairs $(T_{i},\tilde{S}i)|{i=1,...,L}$. Each teaching pair can then be performed m2mKD in parallel. For the $i$-th teaching pair, we replace the $i$-th meta layer with $T_{i}$ and the stitched student module $\tilde{S}_{i}$, giving rise to two hybrid networks:
@@
The two hybrid networks are now considered as“complete” models. Consequently, we can directly apply the conventional response-based knowledge distillation technique. Specifically, for the classification problems considered in this paper, we compare the output logits $z_{T}$ and $z_{\tilde{S}}$ given an input $x$ by measuring the Kullback-Leibler (KL) divergence (Hinton et al.,, 2015). The total loss $\mathcal{L}$ is then defined as the weighted sum of the classification cross-entropy loss $L_{CE}$ and the knowledge distillation loss $L_{KD}$:
@@
End-to-end training. Given a student model $\mathcal{S}$ consisting of $L$ modules, we can run m2mKD for all teaching pairs $(T_{i},\tilde{S}i)|{i=1,...,L}$ in parallel to obtain $L$ distilled student modules $\tilde{S}_{i}^{*}$. The last step is to simply load the learned parameters into $\mathcal{S}$ and perform end-to-end (E2E) training. Note that all stitch layers of the student modules are discarded at this stage. For the NAC model architecture, there can be a dimension mismatch problem when loading the learned parameters for certain components, especially if the datasets used in m2mKD and E2E training differ (refer to Section 4.2). In such cases, these incompatible elements will not be loaded.
@@ 4 Experiments on NACs
NAC is a novel modular architecture as depicted in Figure 3(a). It comprises a read-in layer, multiple propagator layers, and a read-out layer. Modules within two consecutive layers are connected sparsely using Stochastic Kernel Modulated Dot-Product Attention (SKMDPA). The basic computing unit within each module is the ModFC layer, which forms ModFFNs. The majority of parameters are shared across modules, and each module conditions its computation using its own code vector $c_{u}$. The SKMDPA mechanism employs sparse attention to calculate the similarity between modules based on their signature vector $s_{u}$, thereby determining the communication between modules. Each module maintains a state vector $\theta_{u}$, which serves as the input for the subsequent layer. The state vector is initialized in the read-in layer, updated through multiple propagator layers, and eventually used as the input for the read-out layer.

### v2

@@ Abstract
Modular neural architectures are gaining attention for their powerful generalization and efficient adaptation to new domains. However, training these models poses challenges due to optimization difficulties arising from intrinsic sparse connectivity. Leveraging knowledge from monolithic models through techniques like knowledge distillation can facilitate training and enable integration of diverse knowledge. Nevertheless, conventional knowledge distillation approaches are not tailored to modular models and struggle with unique architectures and enormous parameter counts. Motivated by these challenges, we propose module-to-module knowledge distillation (m2mKD) for transferring knowledge between modules. m2mKD combines teacher modules of a pretrained monolithic model and student modules of a modular model with a shared meta model respectively to encourage the student module to mimic the behaviour of the teacher module. We evaluate m2mKD on two modular neural architectures: Neural Attentive Circuits (NACs) and Vision Mixture-of-Experts (V-MoE). Applying m2mKD to NACs yields significant improvements in IID accuracy on Tiny-ImageNet (up to 5.6%) and OOD robustness on Tiny-ImageNet-R (up to 4.2%). Additionally, the V-MoE-Base model trained with m2mKD achieves 3.5% higher accuracy than end-to-end training on ImageNet-1k. Code is available at https://github.com/kamanphoebe/m2mKD.
Keywords:
Knowledge distillation Modular neural architectures
@@ 1 Introduction
Despite the success of large monolithic models in various domains, concerns have emerged regarding their limited generalization ability and increasing computational costs. Meanwhile, modular models have gained attention as promising alternatives to mitigate these drawbacks. In contrast to monolithic models with fixed computational graphs and parameters, modular neural architectures dynamically adapt their parameters to the input, offering favorable properties that are absent in static monolithic models [8]. Unlike monolithic models that optimize parameters collectively, modular models consist of independent modules that can be updated locally without affecting other network parts. These modules specialize in specific tasks and improve generalization performance by activating relevant modules for each input, even for out-of-distribution (OOD) samples [27]. For instance, DEMix Layers [7] jointly represent COVID-19-related data using medical and news modules. Moreover, modular models enhance computational efficiency through conditional calculation. Mixture-of-Experts (MoE), a typical modular architecture [31], significantly increases model capacity while maintaining similar computational requirements to the original model [15, 20].
@@
Although modular architectures surpass monolithic models in terms of OOD robustness and computational efficiency, there is still vast room for improvement in their algorithms and implementations. Training modular models presents challenges due to optimization difficulties arising from sparse interactions. While recent works [42, 24, 26] have investigated the training instability of modular models, thorough studies in this area still need to be completed. On the engineering side, several implementations of MoE exist, but many of them [9, 20] do not consider dynamic workload allocation for experts or support acceleration techniques like mixed precision training. Although adaptive parallelism [14] has addressed the dynamic nature of MoEs, training MoEs remains considerably slower than their monolithic counterparts due to unavoidable communication overheads.
@@
To alleviate the optimization issue, employing a pretrained monolithic model to guide the training of modular models shows potential. Knowledge distillation (KD) [11] is an existing technique that transfers knowledge from a pretrained teacher model to a smaller student model. KD has proven effective in the context of monolithic models. However, directly applying conventional KD approaches to modular models is computationally expensive due to their large model sizes. Using a monolithic model as the teacher for a larger modular model may even harm performance (see Tab. 3). Furthermore, monolithic models trained by regular methods may not be the optimal choice for teachers.
@@
Inspired by the divide-and-conquer training mechanism of Deep Incubation [25], we introduce module-to-module knowledge distillation (m2mKD) to transfer knowledge between sub-modules of the monolithic teacher and the modular student. As illustrated in Figure Fig. 1, we first adopt Deep Incubation, a modular training method, to incubate the teacher modules using a small meta-model. Next, we encourage the student module to imitate the behaviour of the teacher module. Finally, the distilled student modules are used to initialize the modular student model. By adapting the distillation process at the distributed module level, m2mKD significantly reduces the capacity requirement of the teacher model and enables independent training, making distillation less computationally expensive. The cost savings become more noticeable when the modular student is large or when multiple teachers are involved, such as in ensemble learning. Moreover, the teacher model is trained in a modular way, which may benefit teaching a modular student. Note that our m2mKD algorithm does not impose any restrictions on the architecture of both the teacher and student models. We evaluate the performance of m2mKD on both NACs and V-MoE. Experiments show that a NAC model trained using m2mKD improves IID accuracy by approximately 5% on Tiny-ImageNet and enhances out-of-distribution (OOD) robustness by about 4% on Tiny-ImageNet-R. Additionally, there is an average gain of approximately 1% on ImageNet [3] and ImageNet-R [10]. The experimental results for V-MoE models indicate that m2mKD also works in the case of a small teacher.
@@
Our contributions are as follows: (1) To the best of our knowledge, we are the first to investigate knowledge distillation for modular models. (2) We demonstrate the challenges associated with distilling modular models and introduce a tailored algorithm to address these challenges. (3) Our approach can be seen as a promising framework for transforming a monolithic model into a modular model with arbitrary neural architecture. Notably, this transformation is developed in a modular-style manner. (4) The proposed method is capable of handling irregular distillation scenarios where the student model size is larger than the teacher model size. It has the potential to work not only for the monolithic-to-modular case but also for the monolithic-to-monolithic case. (5) We verify the feasibility of using Deep Incubation for modular models.
@@ 2 Related Work
Knowledge distillation. Knowledge distillation is a model compression technique that transfers knowledge from a large teacher network to a smaller student model [11]. Common knowledge distillation approaches involve mimicking the teacher model’s softened output [11, 16, 1]. In scenarios with multiple teachers, the outputs of all teacher models can be averaged [11], or other methods can be used [5, 19, 39]. Traditional knowledge distillation methods often struggle when a significant capacity gap exists between the student and the teacher models. Recent approaches have attempted to address this gap by using teacher assistants [23, 33]. In addition, some existing works conduct distillation at the module level instead of distilling entire models [37, 41, 22]. While previous works primarily focus on compressing model sizes within the context of monolithic models, we propose a module-to-module distillation technique to enhance the performance of modular architectures, especially in unconventional distillation scenarios. Particularly, the distillation of modules is mutually independent and can be executed in a distributed manner.
@@
Modularization. Modular deep learning involves decomposing neural architectures into independent and parameter-efficient modules, where samples are conditionally routed to subsets of modules, and their outputs are aggregated. Modularization has been widely applied in various areas such as transfer learning [21, 12, 28, 13], modular training [25], and scaling up model size [15, 20].
@@
In transfer learning, pre-trained models are often used as modules in an assembled model for new tasks. This assembled model can adapt to new data by adjusting or adding modules to enhance performance in different scenarios. For instance, Brown et al. [2] added a prompt module to the input of a pre-trained model, while Rebuffi et al. [29] introduced adapter modules into the model architecture. Modular training approaches like Deep Incubation [25] incubate modules in individual nodes to avoid communication overhead and accelerate convergence. Conditional computation enables the scaling of model size while maintaining inference complexity. V-MoE [15] scales up vision models with only half the computation required during inference. The MoE framework introduces experts to enable modularization at the FFN layer. Other modular model architectures include NACs [36], as well as [32, 6].
@@
Monolithic to modular. There are several efforts to convert monolithic models into modular architectures. MoEfication [40] directly splits the FFN layers of a monolithic model into multiple experts to form MoE layers, while Sparse Upcycling [17] copies the MLP parameters to the corresponding experts in the MoE layers. In contrast to these works, our proposed m2mKD does not make assumptions about the model architecture and can be applied to any model.
@@ 3 Method
In this section, we first review the essential aspects of Deep Incubation, as our work heavily relies on it. Subsequently, we introduce our proposed method. For convenience, we provide a table (Appendix A) that lists the notations commonly used in this paper.
Figure 2: Overview of Deep Incubation. First, a small meta model is pretrained. Then, the module replaces the corresponding layer in the meta model, and only the module’s parameters are updated during incubation. Finally, the incubated modules are assembled and fine-tuned to obtain the final model.
@@ 3.1 Preliminary: Deep Incubation
Deep Incubation [25] employs a divide-and-conquer strategy to train large models modularly, as illustrated in Fig. 2. The process consists of three stages: meta-model pre-training, module incubation, and assembly. In the initial stage, a small model is pre-trained as a meta model using an end-to-end training approach. Subsequently, each module replaces the corresponding layer in the meta-model, and only the module parameters are updated during the incubation process. The incubation of modules is mutually independent and can be completed in a distributed manner. Finally, the incubated modules are assembled and fine-tuned to obtain the final model. It is worth noting that the number of layers in the meta model corresponds to the number of modules incubated during the incubation stage.
@@ 3.2 Pipeline
Our method follows a pipeline consisting of three steps (Fig. 3): preparation, module-to-module knowledge distillation (m2mKD), and end-to-end (E2E) training. More details are presented below.
@@
As proposed by [25], the initial step involves training the meta model in an end-to-end fashion to incubate the $L$ sub-modules $T_{1},T_{2},...,T_{L}$. Subsequently, these resulting sub-modules are reassembled to form the teacher model $\mathcal{T}$, which undergoes fine-tuning. While we follow the same pipeline described in the original paper, our focus is on the sub-modules themselves rather than the entire model. We opt for the Incubation approach instead of merely separating a pre-trained model because we claim that the module incubation phase imparts additional knowledge to the sub-modules. This allows them to learn how to function as individual modules rather than being incomplete fragments of a whole model.
Module-to-module knowledge distillation. Once the fine-tuning of the assembled model $\mathcal{T}$ is complete, the sub-modules, or we call “teacher modules" hereafter, are ready for running m2mKD. Unlike conventional knowledge distillation approaches, m2mKD aims to transfer knowledge between modules rather than entire models. Similar to the module incubation proposed by [25], we separately link the teacher and student module to the meta model. By comparing the outputs of the two resulting hybrid models, the student modules are encouraged to mimic the behaviour of the corresponding teacher modules. Previous research by Yang et al. [38] demonstrates that blocks located at similar depths in different networks can be considered functionally equivalent. This insight suggests that neural networks learn similar patterns at similar network stages. Exploiting this insight, we assign teacher modules to students at the same depth, resulting in $L$ teaching pairs $(T_{i},\tilde{S}i)|{i=1,...,L}$. Each teaching pair can then be performed m2mKD in parallel. For the $i$-th teaching pair, we replace the $i$-th meta layer with $T_{i}$ and the stitched student module $\tilde{S}_{i}$, giving rise to two hybrid networks:
@@
The two hybrid networks are now considered as “complete" models. Consequently, we can directly apply the conventional response-based knowledge distillation technique. Specifically, for the classification problems considered in this paper, we compare the output logits $z_{T}$ and $z_{\tilde{S}}$ given an input $x$ by measuring the Kullback-Leibler (KL) divergence hinton2015distilling. The total loss $\mathcal{L}$ is then defined as the weighted sum of the classification cross-entropy loss $L_{CE}$ and the knowledge distillation loss $L_{KD}$:
@@
End-to-end training. Given a student model $\mathcal{S}$ consisting of $L$ modules, we can run m2mKD for all teaching pairs $(T_{i},\tilde{S}i)|{i=1,...,L}$ in parallel to obtain $L$ distilled student modules $\tilde{S}_{i}^{*}$. The last step is to simply load the learned parameters into $\mathcal{S}$ and perform end-to-end (E2E) training. Note that all stitch layers of the student modules are discarded at this stage. For the NAC model architecture, there can be a dimension mismatch problem when loading the learned parameters for certain components, especially if the datasets used in m2mKD and E2E training differ (refer to Sec. 4.2). In such cases, these incompatible elements will not be loaded.
@@ 4 Experiments on NACs
NAC is a novel modular architecture as depicted in Fig. 4(a). It comprises a read-in layer, multiple propagator layers, and a read-out layer. Modules within two consecutive layers are connected sparsely using Stochastic Kernel Modulated Dot-Product Attention (SKMDPA). The basic computing unit within each module is the ModFC layer, which forms ModFFNs. The majority of parameters are shared across modules, and each module conditions its computation using its own code vector $c_{u}$. The SKMDPA mechanism employs sparse attention to calculate the similarity between modules based on their signature vector $s_{u}$, thereby determining the communication between modules. Each module maintains a state vector $\theta_{u}$, which serves as the input for the subsequent layer. The state vector is initialized in the read-in layer, updated through multiple propagator layers, and eventually used as the input for the read-out layer.